# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ Acknowledgments
Figure S4: Sources of non-linearity. a) Ideal circuit (no source of non-linearity) layout featuring the memory and buffer modes only. b) SQUID imbalance directly appears in the effective circuit found by Foster synthesis of the environment seen by the SQUID. c) Series inductance in the arms of the SQUID enable a two-Cooper pair tunneling mechanism [47] (double-cross junctions) which is not suppressed at the circuit bias point. d) A superinductor made of finite-length chains of Josephson junctions may be viewed as an inductor in parallel of a chain of inductive elements with quartic potential (spider symbols following the convention of Ref. [41]). In general, the effective modes may present a different impedance $\tilde{Z}_{a,b}$ to this element from the impedance presented to the SQUID. However, given the circuit symmetry, we find $\tilde{Z}_{a,b}=Z_{a,b}$. e) Asymmetric pump waves yield non-linear contributions to the circuit Hamiltonian in the RWA.
@@
S2.1 Asymmetry of the SQUID junctions
We consider the effective two-mode circuit pictured in Fig. S4b, which is obtained after Foster decomposition of the environment seen by the SQUID in the physical circuit (Fig. S4a), and neglecting all modes but the memory and buffer. We derive the SQUID inductive potential when the junctions forming its SQUID are imbalanced. Following the convention of Fig. S4b, we let $\varphi$ the reduced phase across the modes in series with the SQUID, $\Phi_{1}$, $\Phi_{2}$ the fluxes biasing each loop, $E_{J_{1}}$, $E_{J_{2}}$ the energies of each junction. Defining
$\begin{split}\varphi_{\Sigma}&=\frac{\Phi_{1}+\Phi_{2}}{2\varphi_{0}}\\ \varphi_{\Delta}&=\frac{\Phi_{1}-\Phi_{2}}{2\varphi_{0}}\\ E_{\Sigma}&=E_{J_{1}}+E_{J_{2}}\\ E_{\Delta}&=E_{J_{1}}-E_{J_{2}}\end{split}$ (S7)
@@
$V^{{SQUID}}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\varphi-\varphi_{\Delta})$ (S8)
@@
${\bf H}_{asym}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}-E_{\Delta}\cos(\bm{\varphi})$ (S9)
@@
where $i=a,b$ (${\bf a}_{i}$ denoting the annihilation operator of mode $i$), $\bm{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and $\varphi^{ZPF}_{i}=\big{(}\pi Z_{i}/R_{q}\big{)}^{1/2}$. The cosine term in this Hamiltonian features non-linear, photon-number preserving terms that survive the RWA. At leading order (truncating the cosine at fourth order), we find that each mode inherits a self-Kerr non-linearity of amplitude $\chi_{aa}=\frac{1}{2}E_{\Delta}{\varphi_{a}^{ZPF}}^{4}$, $\chi_{bb}=\frac{1}{2}E_{\Delta}{\varphi_{b}^{ZPF}}^{4}$ and both modes are dispersively coupled through a cross-Kerr term $\chi_{ab}=E_{\Delta}{\varphi_{a}^{ZPF}}^{2}{\varphi_{b}^{ZPF}}^{2}$. Given the near perfect symmetry of the ATS in our experiment, these terms are negligible ($\chi_{aa},\chi_{bb},\chi_{ab}<2\pi\times$10 kHz).
S2.2 Stray inductances and multi-Cooper pair tunneling
Figure S5: Inductive dipole made of a junction in series with an inductor. Using Kirchhoff’s law, the dipole potential may be expressed as a function of its generalized phase $\psi$ only.
@@
Following the derivation in Ref. [47], a junction in series with a small inductor (inductive energies $E_{J}\gg\epsilon_{L}$) behaves as an inductive dipole with reduced phase $\psi$ (see Fig. S5). Letting $i$ the current threading the dipole, $\psi_{L}$ and $\psi_{J}$ the reduced phases across the inductor and the junction, we find using Kirchhoff’s law
$\begin{split}\psi_{L}+\psi_{J}=\psi\\ E_{J}\sin(\psi_{J})=\epsilon_{L}\psi_{L}\end{split}$ (S10)
@@
The dipole energy reads $V=-E_{J}\cos(\psi_{J})+\frac{\epsilon_{L}}{2}\psi_{L}^{2}$. Given the hierarchy $E_{J}\ll\epsilon_{L}$, we can solve the equation (S10) to leading order in $E_{J}/\epsilon_{L}$ and inject the solution in the dipole potential to get
$V(\psi)=-E_{J}\cos(\psi)+\frac{E_{J}^{2}}{4\epsilon_{L}}\cos(2\psi)+\mathcal{O}\big{(}\frac{E^{3}_{J}}{\epsilon^{2}_{L}}\big{)}$ (S11)
@@
When replacing standard junctions with two such dipoles in the circuit SQUID (see Fig. S4c), the inductive potential is renormalized as
$\tilde{V}^{{SQUID}}_{\varphi_{\sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\varphi-2\varphi_{\Delta})$ (S12)
where $e_{\Sigma}=\frac{E^{2}_{J}}{2\epsilon_{L}}$ and we have assumed the SQUID to be balanced for simplicity. Now focusing on the ATS operating point at $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,-\pi/2)$, the static circuit Hamiltonian (no pump applied here) reads
@@
${\bf H}_{2\varphi}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}+e_{\Sigma}\cos(2\bm{\varphi})$ (S13)
This term in $\cos(2\bm{\varphi})$ is the dominant contributor to the modes non-linearity at the saddle point. In particular, the memory anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$ (see Fig. 2b). This last property is a consequence of the large zero point fluctuations of the phase of the mode, which are effectively doubled by the Cooper pair pairing mechanism. Indeed, the $\cos(2\bm{\varphi})$ term cannot be truncated at fourth order, which would have yielded a Kerr non-linearity (constant anharmonicity).
@@
We fit the memory energy spectrum with this model (including corrections from the Josephson junctions asymmetry and a non-linear contribution from the circuit superinductors, see following section), $e_{\Sigma}$ being the only free fit parameter. We thus estimate $e_{\Sigma}$ and finally $\epsilon_{L}=h\times 1920~{}$GHz. This inductive energy corresponds to a series inductance of 85 pH in each arm of the SQUID, which matches a rough estimate of the geometric inductance of the tracks (each arm is 120 $\mu$m long, see Fig. S1). We note that a recent paper [69] reported deviations from the standard $\cos(\varphi)$ potential for junctions embedded in several transmon circuits, with significant higher-order harmonics. This effect could contribute to the $\cos(2\varphi)$ term observed in our experiment. We can nevertheless place a conservative upper bound at 0.2 % for the ratio of the second to first harmonic in the junction potential (corresponding to the case where the $\cos(2\varphi)$ term is entirely attributed to intrinsic distortion of the junction potential). This bound is an order of magnitude smaller than the values reported in [69]. We note that our Josephson junctions are fabricated with a long, high-pressure oxidation step (see Sec. S1.2), probably resulting in an oxide barrier thicker than for the devices analyzed in that work. We can thus conclude, in line with the model presented in [69], that junctions with a thick oxide barrier are less prone to the formation of pinholes and are described by a nearly pure $\cos(\varphi)$ potential.
S2.3 Non-linearity induced by the superinductors

### v3

@@ Appendix B Non-idealities and non-linearity of the modes
Figure 8: Sources of non-linearity. a) Ideal circuit (no source of non-linearity) layout featuring the memory and buffer modes only. b) SQUID imbalance directly appears in the effective circuit found by Foster synthesis of the environment seen by the SQUID. c) Series inductance in the arms of the SQUID enable a two-Cooper pair tunneling mechanism [65] (double-cross junctions) which is not suppressed at the circuit bias point. d) A superinductor made of finite-length chains of Josephson junctions may be viewed as an inductor in parallel of a chain of inductive elements with quartic potential (spider symbols following the convention of Ref. [43]). In general, the effective modes may present a different impedance $\tilde{Z}_{a,b}$ to this element from the impedance presented to the SQUID. However, given the circuit symmetry, we find $\tilde{Z}_{a,b}=Z_{a,b}$. e) Asymmetric pump waves yield non-linear contributions to the circuit Hamiltonian in the RWA.
@@
B.1 Asymmetry of the SQUID junctions
We consider the effective two-mode circuit pictured in Fig. 8b, which is obtained after Foster decomposition of the environment seen by the SQUID in the physical circuit (Fig. 8a), and neglecting all modes but the memory and buffer. We derive the SQUID inductive potential when the junctions forming its SQUID are imbalanced. Following the convention of Fig. 8b, we let $\varphi$ the reduced phase across the modes in series with the SQUID, $\Phi_{1}$, $\Phi_{2}$ the fluxes biasing each loop, $E_{J_{1}}$, $E_{J_{2}}$ the energies of each junction. Defining
$\begin{split}\varphi_{\Sigma}&=\frac{\Phi_{1}+\Phi_{2}}{2\varphi_{0}}\\ \varphi_{\Delta}&=\frac{\Phi_{1}-\Phi_{2}}{2\varphi_{0}}\\ E_{\Sigma}&=E_{J_{1}}+E_{J_{2}}\\ E_{\Delta}&=E_{J_{1}}-E_{J_{2}}\end{split}$ (12)
@@ B.1 Asymmetry of the SQUID junctions
$\begin{split}V^{SQUID}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=&-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})\\ &-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\varphi-\varphi_{\Delta})\end{split}$ (13)
@@
${\bf H}_{asym}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}-E_{\Delta}\cos(\boldsymbol{\varphi})$ (14)
@@
where $i=a,b$ (${\bf a}_{i}$ denoting the annihilation operator of mode $i$), $\boldsymbol{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and $\varphi^{ZPF}_{i}=\big(\pi Z_{i}/R_{q}\big)^{1/2}$. The cosine term in this Hamiltonian features non-linear, photon-number preserving terms that survive the RWA. At leading order (truncating the cosine at fourth order), we find that each mode inherits a self-Kerr non-linearity of amplitude $\chi_{aa}=\frac{1}{2}E_{\Delta}{\varphi_{a}^{ZPF}}^{4}$, $\chi_{bb}=\frac{1}{2}E_{\Delta}{\varphi_{b}^{ZPF}}^{4}$ and both modes are dispersively coupled through a cross-Kerr term $\chi_{ab}=E_{\Delta}{\varphi_{a}^{ZPF}}^{2}{\varphi_{b}^{ZPF}}^{2}$. Given the near perfect symmetry of the ATS in our experiment, these terms are negligible ($\chi_{aa},\chi_{bb},\chi_{ab}<2\pi\times$10 kHz).
B.2 Stray inductances and multi-Cooper pair tunneling
[…]
In the first model, each junction and its series inductor is considered as a non-linear inductive element (see Fig. 9a). The circuit is analyzed through Foster synthesis of the linear portion of the circuit connected to these elements (see Fig. 9b). This is the point of view adopted throughout this paper, and it allows us to predict the non-linearity induced on the circuit modes as a function of $\epsilon_{L}$.
In the second model, we analyze the circuit through Foster synthesis of the linear circuit including the series inductors and directly connected to the junctions (see Fig. 9c). This model allows us to extract the value of $\epsilon_{L}$ through finite-element electromagnetic simulation of the circuit.
Figure 9: a) Inductive dipole made of a junction in series with an inductor. Using Kirchhoff’s law, the dipole potential may be expressed as a function of its generalized phase $\psi$ only. b) Circuit partition considered in the model 1 and throughout the paper. We perform Foster synthesis of the linear portion of the circuit colored in gray, yielding a set of normal modes $\{{\bf a}_{i}\}$. This synthesized environment is connected through a single port with phase $\varphi$ (black circles) to a non linear dipole colored in purple (the SQUID, which includes the series inductors accounting for the inductance of tracks and the externally applied magnetic fluxes). The inductive potential of the dipole is expressed as a function of $\varphi$ and of the applied magnetic fluxes $\Phi_{1}$ and $\Phi_{2}$ in Eq. (17). c) Circuit partition considered in the model 2. We synthesize the linear portion of the circuit including the series inductors in the SQUID arms, yielding a set of normal modes $\{{\bf b}_{i}\}$. This environment is connected to the two junctions via 2 ports with with phases $\varphi_{1}$ and $\varphi_{2}$. The potential of each junction is expressed as a function of each applied magnetic flux in Eq. (21). Note that we choose a gauge in which the magnetic fluxes are allocated to the junctions and not to the linear circuit. d) At a DC flux point $\varphi_{\Sigma},\varphi_{\Delta}=(\pi/2,\theta)$, the circuit in (c) is well approximated by a linearized version in which junctions have been replaced with inductances $\pm\frac{L_{J}}{\mathcal{B}\sin(\theta)}.$ This linear circuit is fed to an electromagnetic solver to estimate the inductance of the SQUID arms (see text).
B.2.1 Model 1: dressed Josephson junction
@@
Following the derivation in Ref. [65], a junction in series with a small inductor (inductive energies $E_{J}\gg\epsilon_{L}$) behaves as an inductive dipole with reduced phase $\psi$ (see Fig. 9a). Letting $i$ the current threading the dipole, $\psi_{L}$ and $\psi_{J}$ the reduced phases across the inductor and the junction, we find using Kirchhoff’s law
$\begin{split}\psi_{L}+\psi_{J}=\psi\\ E_{J}\sin(\psi_{J})=\epsilon_{L}\psi_{L}\end{split}$ (15)
@@
The dipole energy reads $V=-E_{J}\cos(\psi_{J})+\frac{\epsilon_{L}}{2}\psi_{L}^{2}$. Given the hierarchy $E_{J}\ll\epsilon_{L}$, we can solve the equation (15) to leading order in $E_{J}/\epsilon_{L}$ and inject the solution in the dipole potential to get
$V(\psi)=-E_{J}\cos(\psi)+\frac{E_{J}^{2}}{4\epsilon_{L}}\cos(2\psi)+\mathcal{O}\big(\frac{E^{3}_{J}}{\epsilon^{2}_{L}}\big)$ (16)
@@
When replacing standard junctions with two such dipoles in the circuit SQUID (see Fig. 8c), the inductive potential is renormalized as
$\begin{split}\tilde{V}^{SQUID}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=&-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})\\ &+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\varphi-2\varphi_{\Delta})\end{split}$ (17)
where $e_{\Sigma}=\frac{E^{2}_{J}}{2\epsilon_{L}}$ and we recall the definition of the reduced magnetic fluxes $\varphi_{\Sigma/\Delta}=\frac{\Phi_{1}~+/-~\Phi_{2}}{2\varphi_{0}}$. Note that we have assumed the SQUID to be balanced for simplicity.
In order to relate predictions from this model to those from the model described below, we note that at a DC flux bias of $(\varphi_{\Sigma},\varphi_{\Delta})=(\pi/2,\theta)$, the static circuit Hamiltonian (no pump applied here) reads
${\bf H}_{\theta}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}-e_{\Sigma}\cos(2\boldsymbol{\varphi}-2\theta)$ (18)
where we remind that $\boldsymbol{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and the $a_{i}$’s are modes found by through Foster synthesis of the linear circuit connected to the two dipoles each formed by a Josephson junction in series with an inductor (see Fig. 9b).
Given that $e_{\Sigma}\ll\omega_{i}$, we can estimate the frequency shift of each mode induced by the second term in the RWA. We first note that, neglecting rotating terms and for low-energy states, $e_{\Sigma}\cos(2\boldsymbol{\varphi}-2\theta)$ is well approximated by $2e_{\Sigma}\mathcal{A}^{4}\cos(2\theta)\boldsymbol{\varphi}^{2}$, where $\mathcal{A}=e^{-\frac{1}{2}\sum_{i}{\varphi_{i}^{ZPF}}^{2}}$ and $\mathcal{A}^{4}$ is a prefactor appearing when normal-ordering the operator $\cos(2\boldsymbol{\varphi})$. Still in the RWA, this term provides each mode with an additional inductive energy that adds to the energy $E_{L_{i}}=\hbar\omega_{i}{\varphi_{i}^{ZPF}}^{2}/2$ determined by the Foster synthesis of the linear circuit. Compared to the case where the series inductors in the SQUID are neglected, we find that the modes’ resonance frequencies to be dressed as
$\omega_{i}\rightarrow\omega_{i}\Big(1+\frac{\mathcal{A}^{4}E_{J}^{2}}{\epsilon_{L}E_{L_{i}}}\cos(2\theta)\Big)$ (19)
B.2.2 Model 2: dressed modes
We now adopt a different viewpoint and synthesize the linear portion of the circuit including the inductors in the SQUID arms as depicted in Fig. 9c. The synthesis yields a set of modes $\{{\bf b}_{i}\}$ and allows us to express the phase of each mode across the two ports to which the Josephson junctions are connected:
$\boldsymbol{\varphi}_{1}=\sum_{i}\varphi^{ZPF}_{i,1}({\bf b}_{i}+{\bf b}_{i}^{\dagger})\qquad\qquad\boldsymbol{\varphi}_{2}=\sum_{i}\varphi^{ZPF}_{i,2}({\bf b}_{i}+{\bf b}_{i}^{\dagger})$ (20)
Given the circuit symmetry, each mode either has symmetric phases across the two ports ($\varphi^{ZPF}_{i,2}=\varphi^{ZPF}_{i,1}$), or antisymmetric phases ($\varphi^{ZPF}_{i,2}=-\varphi^{ZPF}_{i,1}$). We now label with an index $k_{\Sigma}$ the modes with symmetric phases (with phase vacuum fluctuations $\varphi^{ZPF}_{k_{\Sigma}}$), and with an index $k_{\Delta}$ the modes with antisymmetric phases (with vacuum fluctuations $\varphi^{ZPF}_{k_{\Delta}}$).
In a similar spirit to the analysis made in this paper introduction, we can write the inductive potential of each junction as
$\begin{split}&V_{1}(\varphi_{1})=-E_{J}\cos(\varphi_{1}-\varphi_{\Sigma}-\varphi_{\Delta})\\ &V_{2}(\varphi_{2})=-E_{J}\cos(\varphi_{2}+\varphi_{\Sigma}-\varphi_{\Delta})\end{split}$ (21)
where we have chosen a gauge in which the externally applied magnetic fluxes are allocated to the junctions. Expressing $\boldsymbol{\varphi}_{1}$ and $\boldsymbol{\varphi}_{2}$ in the normal mode basis, the circuit Hamiltonian at the DC flux point $(\varphi_{\Sigma},\varphi_{\Delta})=(\pi/2,\theta)$
@@ B.2.2 Model 2: dressed modes
$\tilde{{\bf H}}_{\theta}=\sum_{k_{\Sigma}}\hbar\omega_{k_{\Sigma}}{\bf b}^{\dagger}_{k_{\Sigma}}{\bf b}_{k_{\Sigma}}+\sum_{k_{\Delta}}\hbar\omega_{k_{\Delta}}{\bf b}^{\dagger}_{k_{\Delta}}{\bf b}_{k_{\Delta}}-2E_{J}\sin\Big(\sum_{k_{\Delta}}\varphi^{ZPF}_{k_{\Delta}}({\bf b}_{k_{\Delta}}+{\bf b}^{\dagger}_{k_{\Delta}})\Big)\cos\Big(\sum_{k_{\Sigma}}\varphi^{ZPF}_{k_{\Sigma}}({\bf b}_{k_{\Sigma}}+{\bf b}^{\dagger}_{k_{\Sigma}})-\theta\Big)$ (22)
As we are concerned with predicting the modes’ frequency dispersion as a function of $\theta$ and in the low-energy regime, we now approximate the Josephson term to leading order and drop linear drive terms that do not impact the modes’ frequencies
$\tilde{{\bf H}}_{\theta}\approx\sum_{k_{\Sigma}}\hbar\omega_{k_{\Sigma}}{\bf b}^{\dagger}_{k_{\Sigma}}{\bf b}_{k_{\Sigma}}+\sum_{k_{\Delta}}\hbar\omega_{k_{\Delta}}{\bf b}^{\dagger}_{k_{\Delta}}{\bf b}_{k_{\Delta}}-2\mathcal{B}E_{J}\sin(\theta)\Big(\sum_{k_{\Delta}}\varphi^{ZPF}_{k_{\Delta}}({\bf b}_{k_{\Delta}}+{\bf b}^{\dagger}_{k_{\Delta}})\Big)\Big(\sum_{k_{\Sigma}}\varphi^{ZPF}_{k_{\Sigma}}({\bf b}_{k_{\Sigma}}+{\bf b}^{\dagger}_{k_{\Sigma}})\Big)$ (23)
where the prefactor $\mathcal{B}=e^{-\frac{1}{2}\sum_{k_{\Sigma}}{\varphi_{k_{\Sigma}}^{ZPF}}^{2}-\frac{1}{2}\sum_{k_{\Delta}}{\varphi_{k_{\Delta}}^{ZPF}}^{2}}$ appears when normal ordering the non-linear operators. Finally, we note that this Hamiltonian encodes the dynamics of the linear circuit depicted in Fig. 9d, in which the Josephson junctions are replaced with linear inductors with inductances $\frac{L_{J}}{\mathcal{B}\sin(\theta)}$ and $\frac{L_{J}}{\mathcal{B}\sin(\theta)}$.
Before moving on to estimating the inductance of the tracks of the SQUID, we make two observations
  • We kept our formalism general and indexed modes with labels $k_{\Sigma}$, $k_{\Delta}$. Indeed, in a realistic setting including the circuit feedlines terminated by a 50 $\Omega$ environment, the number of modes to include in the Foster synthesis is infinite. Nevertheless, among symmetric modes, the synthesis is dominated by three modes that can be identified to the memory, buffer and readout modes. Other modes are expected to resonate at much higher frequency and to participate only weakly in the junctions. On the anti-symmetric side, the modes have a low impedance or do not participate strongly in the ATS (only a mode corresponding to charge oscillations between the intrinsic junction capacitors is expected to participate strongly but resonates at a much higher frequency than the modes considered so far and is of low impedance). Overall we can approximate the Josephson energy reduction factor $\mathcal{B}$ with the factor $\mathcal{A}$ found in the model 1.
  • In the model 1, the frequency dispersion of the circuit modes as a function of the applied differential flux $\theta$ is the consequence of an additional inductive energy stemming from the (unsuppressed) tunneling of pairs of Cooper pairs through the junction. In the model 2, the frequency dispersion comes from the hybridization of the symmetric modes with the anti-symmetric environment. However, since choosing where to set the limit of the linear portion of the circuit to be synthesized is arbitrary, the two models should yield identical results and therefore these two mechanisms are equivalent. We note that in the model 2, the coupling factor of symmetric modes to anti-symmetric ones depends on $\theta$ as $g\propto\sin(\theta)$ such that in the weak hybridization limit, we expect a frequency dispersion $\propto\sin^{2}(\theta)$, compatible with the dispersion found in the model 1.
B.2.3 Non-linearity of the modes and estimates for the inductance of the SQUID arms.
In principle, one can estimate the inductance $\varphi_{0}^{2}/\epsilon_{L}$ of the SQUID tracks by recording the modes’s resonance frequency as a function of applied DC fluxes $(\varphi_{\Sigma},\varphi_{\Delta})=(\pi/2,\theta)$ and fitting the obtained data with Eq. (19). However, this data was not acquired in our experiment. In the following, we rather estimate $\epsilon_{L}$ through the memory mode anharmonicity pattern at the saddle point. Subsequently, we compare this estimate to the value found in electromagnetic simulations based on the model 2.
Now focusing on the ATS operating point at $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,-\pi/2)$, the static Hamiltonian derived in the model 1 reads
${\bf H}_{2\varphi}=\sum_{i}\hbar\omega_{i}{\bf a}^{\dagger}_{i}{\bf a}_{i}+e_{\Sigma}\cos(2\boldsymbol{\varphi})$ (24)
We fit the memory energy spectrum recorded at the ATS saddle point (see Fig. 2b) with this model, thus estimating $e_{\Sigma}$ and finally $\epsilon_{L}=h\times 1920$~GHz. This inductive energy corresponds to a series inductance of 85 pH in each arm of the SQUID. Note that we include in our fit corrections from the Josephson junctions asymmetry (independently calibrated) and a non-linear contribution from the circuit superinductors (see following section).
As a sanity check, we estimate the geometric inductance of the tracks of the SQUID leveraging the equivalence between model 1 and model 2. In electromagnetic simulations (with HFSS by Ansys), we replace the junctions in our circuit with linear lumped elements with inductance $\pm\frac{L_{J}}{\mathcal{B}\sin(\theta)}$ (see comment below Eq. (23) and Fig. 9d). For consistency with the model 1, we do not include intrinsic junction capacitance in these simulations. For each value of $\theta$, the EM solver returns the dressed frequency of each mode. We then fit the frequency dispersion of each of the three modes with Eq. (19), yielding three estimates for $\epsilon_{L}$. The corresponding inductances are 7979~pH, 8888~pH and 7474~pH respectively, in good agreement with the value estimated from the memory mode anharmonicity. We also tried a more direct approach in which the inductance of the whole ATS (comprising the SQUID and the shunt inductor) is estimated through EM simulations before subtracting the contribution from the junctions and the shunt inductor. This approach yielded a slightly higher estimate of 100 pH for the inductance of the SQUID arms. We consider this direct approach as less reliable as the ATS’s inductance is estimated from the impedance seen between two particular points of the circuit, which does not account for the finite size of the circuit tracks.
@@
Here, we note that a recent paper [83] reported deviations from the standard $\cos(\varphi)$ potential for junctions embedded in several transmon circuits, with significant higher-order harmonics attributed to pinholes in the junction oxide barrier. In our case, the observed $\cos(2\varphi)$ term can be entirely attributed to geometric inductance in the SQUID arms. Moreover, the amplitude of this term is only 0.2 % of the amplitude of the standard $\cos(\varphi)$ contribution. This ratio is an order of magnitude lower than the values reported in [83]. Since our Josephson junctions are fabricated with a long, high-pressure oxidation step (see Sec. A.2), their oxide barrier is probably thicker than for junctions analyzed in that work. We can thus conclude, in line with the model presented in [83], that junctions with a thick oxide barrier are less prone to the formation of pinholes and are described by a nearly pure $\cos(\varphi)$ potential.
To conclude this section, we note that the $\cos(2\boldsymbol{\varphi})$ term is the dominant contributor to the modes non-linearity at the saddle point. In particular, it sets the memory anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$, which is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$ (see Fig. 2b). This last property is a consequence of the large zero point fluctuations of the phase of the mode, which are effectively doubled by the Cooper pair pairing mechanism. Indeed, the $\cos(2\boldsymbol{\varphi})$ term cannot be truncated at fourth order, which would have yielded a Kerr non-linearity (constant anharmonicity). Beyond introducing non-linearity into the memory mode, which prevents stabilizing a full four-legged cat state, we do not expect the stray inductance of the SQUID arms to impact significantly the results and calculations presented in this work. In particular, while series inductors slightly dilute the participation of the Josephson junctions in the circuit modes, the impact on the four-to-one photon exchange rate is negligible (see Table 2: more than 99 % of the modes’s phase fluctuations are supported by the junctions and not by the series inductors).
B.3 Non-linearity induced by the superinductors